OpenStudy (anonymous):

Solve the following equation for p. Please show all of your work. r=4q-5p/9

OpenStudy (saifoo.khan):

9r = 4q - 5p 9r -4q = -5p -9r + 4q = 5p (-9r + 4q)/5 = p